A body starts rest and moves with a uniform accelaration of 2m/s2 in straight line

a. what is its velocity after 5s?

b.how far has it travelled in this time?

c.when will it be 100m from the starting point?


1
Expert's answer
2022-06-01T12:49:42-0400

(a)

"v=v_0+at=0+2*5=10\:\rm m/s"

(b)

"d=v_0t+at^2/2=0*5+2*5^2/2=25\:\rm m"

(c)

"t=\sqrt{\frac{2d_1}{a}}=\sqrt{\frac{2*100}{2}}=10\:\rm s"